OpenStudy (anonymous):

Someone please help!. A student puts two dynamics carts with a speed bumper between them on a track and presses them together. The total mass of the carts is 3.0 kg. Once the student releases them, the carts spring apart and roll away from each other. One cart has a mass of 2.0 kg and a final velocity of 2.5 m/s [S]. Calculate the final velocity of the other cart

OpenStudy (anonymous):

First, we need to calculate the momentum of the system before the carts roll away from one another. Can you do this?

OpenStudy (anonymous):

p = mv momentum = mass * velocity

OpenStudy (anonymous):

yeah, one second

OpenStudy (anonymous):

For the moment I did, p=(3.0kg)(-2.5m/s) =-7.5

OpenStudy (anonymous):

momentum*

OpenStudy (anonymous):

Ok. That's after the carts have pushed away from one another. We'll need that in a moment, though. What's the initial momentum, before they start moving away from each other?

OpenStudy (anonymous):

I'm not sure if this is correct, but wouldn't it be 0 since both they were both at rest before they sprung apart?

OpenStudy (anonymous):

Exactly! Well done! So, they start with zero momentum. We know that momentum is conserved, so after they push away, they'll also have zero total momentum. You've calculated the momentum of one of the carts, already. We can use these facts to find the momentum of the second cart.

OpenStudy (anonymous):

That is: \[0 = P_{cart1}+P{cart2} = -7.5 + P_{cart2}\] Note, you can make the 7.5 positive by redefining your coordinates, if you wish. So, what is P cart2?

OpenStudy (anonymous):

I hope this is right. Sorry I'm not very confident. p=(2.5)(2)=5?

OpenStudy (anonymous):

sorry -5

OpenStudy (anonymous):

Nope. Here's where we are so far: \[P_{initial} = P_{final}\]Conservation of Momentum \[P_{initial} = 0\]Therefore, \[P_{final} = 0\] \[P_{final} = P_{cart1} + P{cart2}\] \[P_{cart1} = -7.5 \frac{kg*m}{s}\] Putting this together: \[0 = -7.5 + P_{cart2}\] So: \[P_{cart2} = 7.5\] Now, both carts are 3kg together. Cart 1 is 2 kg. What is the mass of Cart 2?

OpenStudy (anonymous):

5kg

OpenStudy (anonymous):

3 - 2 = ?

OpenStudy (anonymous):

sorry 1kg

OpenStudy (anonymous):

Yep. So: \[P_{cart2} = 7.5 = m*v\] m = 1, so what is v?

OpenStudy (anonymous):

7.5m/s

OpenStudy (anonymous):

There you go :)

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

My pleasure